Required information

Exercise 13-6 Stock dividends and per share book values LO P2

[The following information applies to the questions displayed below.]

The stockholders’ equity of TVX Company at the beginning of the day on February 5 follows:

Common stock—$15 par value, 150,000 shares
authorized, 56,000 shares issued and outstanding
$ 840,000
Paid-in capital in excess of par value, common stock 525,000
Retained earnings 675,000
Total stockholders’ equity $ 2,040,000


On February 5, the directors declare a 10% stock dividend distributable on February 28 to the February 15 stockholders of record. The stock’s market value is $33 per share on February 5 before the stock dividend. The stock’s market value is $30 per share on February 28.

Exercise 13-6 Part 1

1. Prepare entries to record both the dividend declaration and its distribution.
2. One stockholder owned 950 shares on February 5 before the dividend. Compute the book value per share and total book value of this stockholder’s shares immediately before and after the stock dividend of February 5. (Round your "Book value per share" answers to 3 decimal places.)

3. Compute the total market value of the investor’s shares in part 2 as of February 5 and February 28.
Answer #1

1.

Date General Journal Debit Credit
Feb. 05 Retained earnings (10% x 56000 x $33) 184800
Common stock dividend distributable (10% x 56000 x $15) 84000
Paid-in capital in excess of par value, common stock 100800
(To record declaration of 10% stock dividend)
Feb. 28 Common stock dividend distributable 84000
Common stock 84000
(To record distribution of common stock dividend)

2.

Before After
Book value per share $                                                                                                    36.429 $       33.117
Total book value of shares $                                                                                                    34,608 $       34,607

Note: There is a $1 difference in the total book value of shares before and after the stock dividend due to rounding off.

Working:

Book value per share = Total stockholders' equity/Number of shares of common stock outstanding

Before:

Book value per share = $2040000/56000 = $36.42865 = $36.429

Total book value of shares = 950 x $36.429 = $34607.55 = $34608

After:

The stock dividend will not change the total stockholders' equity though there will be a change intra-accounts within stockholders' equity. The number of shares of common stock outstanding will increase to the tune of the stock dividend, which will result in a decline in the book value per share.

Stock dividend = 10% x 56000 = 5600

Number of shares of common stock outstanding = 56000 + 5600 = 61600

Book value per share = $2040000/61600 = $33.1168 = $33.117

Number of shares post stock dividend = 950 + (10% x 950) = 950 + 95 = 1045

Total book value of shares = 1045 x $33.117 = $34607.27 = $34607

3.

February 5 February 28
Total market value of shares $                                                                                                    31,350 $       31,350
(950 x $33) (1045 x $30)